OpenStudy (anonymous):

simplify each product help? sqrt 6 ( sqrt 2 + sqrt 3)

OpenStudy (anonymous):

i know it suppose to come out as 7.7 but i keep getting 7.4?

OpenStudy (anonymous):

\[\sqrt{6}\left(\sqrt{2}+\sqrt{3}\right)=3 \sqrt{2}+2 \sqrt{3}=4.24264+3.4641=7.70674 \]